After stricking a floor , a certain ball rebounds 4/5 th height from it has fallen. Find the total distance that it travels before coming to rest, if it is gently dropped from a height of 120 m
Answer by stanbon(75887)

After stricking a floor , a certain ball rebounds 4/5 th height from it has fallen. Find the total distance that it travels before coming to rest, if it is gently dropped from a height of 120 m
----
Infinite Geometric Series with a= 120 meters and r= 4/5
----------------------
Total distance = Sum = a[1/(1-r)] = 120[1/(1/5)] = 600 ft
----------------------
Cheers,
Stan H.